Heights of men are normally distributed with a mean of 176 cm and a standard deviation of 7 cm. What is the approximate percentage of men greater than 155 cm?
Answer #1

Solution :

Given that ,

mean = = 176

standard deviation = = 7

P(x > 155) = 1 - p( x< 155)

=1- p P[(x - ) / < (155 - 176) / 7 ]

=1- P(z < -3)

Using z table,

= 1 - 0.0087

= 0.9987

The percentage is = 99.87%